OpenStudy (anonymous):

In my past account, no one was able to solve this. I wonder if someone can solve this now. Prove: \[S_n = \frac n2 (2a + (n-1)d)\]

OpenStudy (anonymous):

no substitution allowed (by that i mean brutely substituting every number to prove it's right) and no mathematical induction either

OpenStudy (anonymous):

Sn being the sum of any arithmetic progression??

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

a1+a2+a3+a4+a5+....+an=a1+(a1+d)+(a1+2d)+......+(a1+(n-1)d) then, using conmutative property Sn=a1+a2+a3+.....+an= (a1+a1+a1+a1+a1+a1+a1+a1+..+a1)+(d+2d+.....(n-1)d) Sn=a1*n+(d+2d+....(n-1)d) factorize by d Sn=n*a1+d(1+2+3+....(n-1)) which is the sum of the first n-1 natural numbers (which is equal to (n-1)n/2 ) Sn=n*a1+d(n(n-1)/2) \[S_n=n(a1+\frac{ d }{ 2 } (n-1))\] then just multiply and divide by 2 on the right side \[S_n=\frac{ n }{ 2 } (2a_1+d(n-1))\] works for you?

OpenStudy (anonymous):

isn't this a variation of mathematical induction?

OpenStudy (anonymous):

but not induction itself xD

OpenStudy (anonymous):

lol. unfair :p well i'll see first if anyone can solve it without mathematical induction

OpenStudy (blockcolder):

Use what Gauss used to solve the 1+2+...+100 problem: Write the sum twice, once forward, once backward.

OpenStudy (anonymous):

oh that thing?

OpenStudy (blockcolder):

To elaborate: \[S=a+(a+d)+(a+2d)+\dots+(a+(n-1)d)\\ S=(a+(n-1)d)+(a+(n-2)d)+\dots+(a+d)+a\] Add them, and notice that there are n terms on RHS. Each of these add up to \(2a+(n-1)d\): \[2S=n(2a+(n-1)d)\] and the result follows.

OpenStudy (anonymous):

but...isn't this to infinity? how do you divide infinity by 2?

OpenStudy (anonymous):

ah. that

OpenStudy (anonymous):

hmm yes. you cross multiplied. then what?

OpenStudy (loser66):

you guys are genius. wish have more medal to give

OpenStudy (blockcolder):

Well, it's not obvious because I don't know how to align things. The first term of the RHS in the last sum is \(a+(a+(n-1)d)=2a+(n-1)d\), the next term is \((a+d)+(a+(n-2)d=2a+(n-1)d\), etc. There are \(n\) of these, in total.

OpenStudy (blockcolder):

Does this work for you?

OpenStudy (texaschic101):

s = a + a+d +.....a + (n - 2)d + a + (n -1)d now write s in reverse order... s = a + (n-1)d + a + (n-2)d Now add them....term by term s = a + a + d + a + (n-2)d + a + (n-1)d s = a + (n-1)d + a + (n-2)d + a + d + a ----------------------------------------------------------- 2s = (2a + (n-1)d + (2a + (n-1)d + (2a + n-1)d......+ (2a + (n-1)d Each term is the same and there are " n " of them so : s = n x (2a + (n-1)d) so just divide by 2 and we get : s = n/2 x (2a (n-1)d)

OpenStudy (texaschic101):

oops forgot something....at the top, in reverse order s = a + (n-1)d + a + (n-2)d + (a + d) + a

OpenStudy (anonymous):

what does s mean in your solution texas?

OpenStudy (anonymous):

is it also sum?

OpenStudy (texaschic101):

yes..it is sum

OpenStudy (anonymous):

i must admit...these three are smarter than 3 green users =))) 3 green users were unable to solve it in the past

OpenStudy (texaschic101):

it can be confusing :)

OpenStudy (blockcolder):

That's what reading The Art and Craft of Problem Solving does to you.

OpenStudy (texaschic101):

so true :)

OpenStudy (anonymous):

i would show you what they answered....but that would mean revealing the identity of my permanently banned account...so i cant...especially in the presence of a mod :S

OpenStudy (texaschic101):

Yes, I agree.....I wouldn't take any chances